# Low-Altitude Wireless Networks: A Comprehensive Survey

### v2

@@ 7.2 Low-Altitude ATM
Apart from effective LAA structuring, the management of LAA has become a critical issue with the rise of UAM. To address the challenges associated with increasing traffic densities, macroscopic fundamental diagrams (MFDs), originally developed for ground traffic management, have been adapted for LAA [Rastgoftar2024]. An MFD characterizes the aggregate relationship among traffic flow, density, and speed at the network level and thereby provides a compact description of the operational state of the airspace. As shown in Fig. 9, the simulation data and the fitted curve jointly reveal a characteristic three-regime behavior. In the free-flow regime, at low densities, the normalized traffic flow increases almost linearly with density. As density approaches a critical value, the flow reaches a maximum, which indicates the practical capacity of the airspace. Beyond this critical density, the system enters a congested regime, where additional aircraft lead to a reduction in flow and the emergence of unstable operating conditions. This relationship demonstrates that MFDs can capture capacity limits and congestion onset in LAA, and thus provide a useful tool for designing density thresholds, scheduling policies, and flow control strategies for LAWNs. The work of [Bulusu2018] presented a throughput-based capacity metric to evaluate the impact of increasing air traffic on LAA. The proposed method simulated traffic and analyzed the effects of different conflict resolution methods, where the findings suggest that throughput decreases as traffic density rises, even while maintaining safety, offering a new way to assess conflict detection strategies. The authors of [Li2022] addressed the challenges faced by the national airspace system in optimizing traffic flow, in which various strategies for managing air traffic are reviewed. The authors emphasized the role of collaborative decision-making (CDM) and automation to improve capacity and reduce congestion in high-density environments. Furthermore, the study of [Work2009] introduced a new approach to air traffic flow optimization using convex formulations of partial differential equations (PDEs). The authors modeled air traffic flow in an Eulerian network and demonstrated how to convert these models into linear and quadratic optimization problems, providing globally optimal solutions for large-scale air traffic systems.
@@
A machine learning-based decision support system for ATM was further developed by integrating explainable AI (XAI) in [aerospace8080224]. By employing learning-based algorithms to predict risks, this approach can assist controllers in managing air traffic effectively. The incorporation of XAI ensures that the reasoning behind AI decisions is transparent and understandable to human operators, thereby increasing trust in automated systems. Furthermore, the research explored the system’s application in both controlled and uncontrolled airspace, illustrating its potential to enhance safety and operational efficiency in low-altitude ATM. To achieve efficient ATM in real-world environments, significant reforms are being made in the strategic planning and development of ATM systems under the single European sky (SES) initiative [REZO20235], motivated by the need to establish unified strategic objectives to improve LAWNs’ capacity, safety, and environmental sustainability. A comprehensive review of the ATM planning framework in Europe has revealed critical gaps, particularly regarding the alignment of performance targets with local operational needs. The findings suggest that current strategic goals for capacity are overly broad and fail to address the specific requirements of individual regions and airspaces. To sum up, although the low-altitude ATM development is progressing, it remains in the early stages to support practical implementation.
@@
Moreover, the deployment of LAWNs necessitates a strategic alignment with both international frameworks and domestic regulations to ensure safety, interoperability, and compliance. A comprehensive review of representative standards relevant to LAWNs is provided to demonstrate how such systems can seamlessly integrate into global and national UAV ecosystems while adhering to regulatory requirements. The ICAO has developed the ATM framework, which outlines essential capabilities for managing UAVs within LAA [icao_utm_framework]. This framework emphasizes the need for global harmonization in airspace management and the integration of UAVs into existing ATM systems. Key capabilities within the ATM system include flight authorization, real-time information sharing, conflict management, and automated traffic de-confliction, all of which must be supported by LAWNs to ensure safe and coordinated UAV operations. By doing so, LAWNs can ensure the safe management of UAVs, complementing the existing aviation infrastructure.
@@
In the United States, the Federal Aviation Administration (FAA) introduced its ATM concept of operations [FAA2022], envisioning a network of UAV service suppliers that provide services such as digital flight planning, airspace management, and UAV tracking. The FAA’s framework operates through cloud-to-cloud data exchanges, eliminating the need for traditional voice communication with air traffic control. A key feature of the FAA system is Remote ID, which mandates UAVs to broadcast unique identification and location data to enable real-time tracking and compliance with airspace restrictions. LAWNs must integrate with these services, ensuring that UAVs are monitored, controlled, and managed in accordance with FAA regulations. To support studies on future parcel delivery by UAV, the Japan Aerospace Exploration Agency (JAEA) has introduced a dedicated simulation platform for the 2030 timeframe, termed the scalable simulator for knowledge of low-altitude environment (SKALE). This framework emulates BVLoS operations above densely populated urban regions. As illustrated in Fig. 10, a representative delivery scenario is configured with a nominal cruise altitude of 120 m. Analysis based on SKALE indicates that current conflict detection and resolution strategies are not yet sufficient for large-scale, highly automated traffic. In particular, the appropriate setting and modelling of separation minima and flight speeds remain open research problems that require further refinement.
Figure 10: The typical UAV delivery workflow of Ref. [oosedo2021unmanned].
@@
Similarly, the Civil Aviation Administration of China (CAAC) has implemented regulations that follow a risk-based approach to managing LAA [caac_low_altitude_service_system_plan_2018]. UAVs are categorized based on their weight and operational risk level, with all UAVs, except the smallest, required to be equipped with unique electronic identification and registered under a real-name system. In addition, LAWNs are required to integrate geofencing capabilities and enforce remote flight restrictions according to the operational limitations set by regulatory authorities. The policies within China are actively driving the LAE development, generating significant demand for high-capacity, scalable LAWNs that can support dense UAV applications in terms of drone delivery, eVTOL air taxis, and remote sensing. To meet these requirements, LAWNs must provide ultra-reliable communication and low-latency connectivity to ensure seamless integration and compliance with regulatory standards in an expanding LAA. Toward this end, the 3rd Generation Partnership Project (3GPP) has released several technical reports (TRs) addressing LAWN communications [3gpp_tr_36777]. Notably, TR 36.777 explores the UAV connectivity by leveraging LTE, highlighting key challenges such as interference and mobility. Furthermore, TR 38.811 and TR 38.821 focus on the integration of 5G new radio (NR) within non-terrestrial networks (NTN), which includes satellites and HAPs.
@@ 7.3 Emerging Path Planning Technology
The preceding discussion on low-altitude ATM focuses on macroscopic aspects such as capacity, separation policy, and regulatory coordination. In practice, these system-level concepts are ultimately realized through the individual trajectories flown by each aerial vehicle. Therefore, path planning forms the microscopic counterpart of ATM and directly determines whether a low-altitude airspace can accommodate dense operations without compromising safety or efficiency. A well-designed trajectory must remain compatible with corridor structures, restricted zones, and dynamic flow constraints, while obeying platform limits on energy, payload, and maneuverability. In the context of LAWNs, where numerous heterogeneous UAVs operate in close proximity, deficiencies in path design manifest as conflict hot spots, unnecessary detours, and excessive delay. From a design perspective, most emerging path planning technologies are driven by two primary objectives, namely, trajectory safety and trajectory efficiency [aggarwal2020path]. Specifically, safety-oriented planning emphasizes collision and obstacle avoidance, while compliance with no-fly regions and separation minima. On the other hand, efficiency-oriented planning seeks to reduce mission cost, for example, by shortening flight time and lowering energy consumption through multiobjective criteria. In many recent studies, safety is enforced through hard constraints, whereas efficiency guides the selection of a preferred solution within the feasible set.
@@
By contrast, efficiency-oriented path planning methods build on this safety foundation and place additional emphasis on operating cost and computational complexity. For instance, the work in [77] modeled UAV path planning under partial observability as a partially observable Markov decision process and sought trajectories that minimize cumulative operating cost while accounting for nonlinear dynamics and measurement uncertainty. This formulation provides a principled way to balance performance and robustness when the vehicle has only partial knowledge of its environment. Along this line, the work in [cui2014improved] observed that traditional UAV path planners often ignore mission-specific constraints such as terminal heading and proposed a bidirectional ant-colony-optimization scheme on a grid-based workspace to address this limitation. Ant colonies are initialized near the start and goal configurations and expand simultaneously along predetermined directions, after which a dedicated path-selection mechanism combines and filters candidate routes. Improved pheromone-update rules and successor-selection strategies were introduced to accelerate convergence and reduce the risk of local optima. From a cooperative perspective, the authors in [belkadi2019design] developed a distributed controller for a fleet of UAVs under a multi-agent-systems framework, where online path planning is formulated as a particle-swarm-optimization (PSO) problem solved independently on each vehicle. By minimizing a suitably designed cost function, each local PSO instance determines a trajectory that contributes to formation keeping, target tracking, and collision avoidance, and experiments with fixed and moving targets, external disturbances, and agent loss confirmed that the distributed strategy can maintain coherent fleet behavior and safety without centralized coordination. Overall, emerging path planning technologies complement low-altitude ATM by bridging system-level policies with executable trajectories. Safety-oriented configuration-space methods and cooperative decision frameworks provide reliable guarantees in structured and moderately dynamic airspace, whereas efficiency-oriented approaches enhance scalability and performance in large-scale, high-density LAWNs. The combined use of these approaches is expected to be essential for supporting dense, task-driven operations while maintaining both operational safety and mission effectiveness.

### v3

@@ 7.2 Low-Altitude ATM
Apart from effective LAA structuring, the management of LAA has become a critical issue with the rise of UAM. To address the challenges associated with increasing traffic densities, macroscopic fundamental diagrams (MFDs), originally developed for ground traffic management, have been adapted for LAA [134]. An MFD characterizes the aggregate relationship among traffic flow, density, and speed at the network level and thereby provides a compact description of the operational state of the airspace. As shown in Fig. 9, the simulation data and the fitted curve jointly reveal a characteristic three-regime behavior. In the free-flow regime, at low densities, the normalized traffic flow increases almost linearly with density. As density approaches a critical value, the flow reaches a maximum, which indicates the practical capacity of the airspace. Beyond this critical density, the system enters a congested regime, where additional aircraft lead to a reduction in flow and the emergence of unstable operating conditions. This relationship demonstrates that MFDs can capture capacity limits and congestion onset in LAA, and thus provide a useful tool for designing density thresholds, scheduling policies, and flow control strategies for LAWNs. The work of [24] presented a throughput-based capacity metric to evaluate the impact of increasing air traffic on LAA. The proposed method simulated traffic and analyzed the effects of different conflict resolution methods, where the findings suggest that throughput decreases as traffic density rises, even while maintaining safety, offering a new way to assess conflict detection strategies. The authors of [89] addressed the challenges faced by the national airspace system in optimizing traffic flow, in which various strategies for managing air traffic are reviewed. The authors emphasized the role of collaborative decision-making (CDM) and automation to improve capacity and reduce congestion in high-density environments. Furthermore, the study of [165] introduced a new approach to air traffic flow optimization using convex formulations of partial differential equations (PDEs). The authors modeled air traffic flow in an Eulerian network and demonstrated how to convert these models into linear and quadratic optimization problems, providing globally optimal solutions for large-scale air traffic systems.
@@
A machine learning-based decision support system for ATM was further developed by integrating explainable AI (XAI) in [175]. By employing learning-based algorithms to predict risks, this approach can assist controllers in managing air traffic effectively. The incorporation of XAI ensures that the reasoning behind AI decisions is transparent and understandable to human operators, thereby increasing trust in automated systems. Furthermore, the research explored the system’s application in both controlled and uncontrolled airspace, illustrating its potential to enhance safety and operational efficiency in low-altitude ATM. To achieve efficient ATM in real-world environments, significant reforms are being made in the strategic planning and development of ATM systems under the single European sky (SES) initiative [135], motivated by the need to establish unified strategic objectives to improve LAWNs’ capacity, safety, and environmental sustainability. A comprehensive review of the ATM planning framework in Europe has revealed critical gaps, particularly regarding the alignment of performance targets with local operational needs. The findings suggest that current strategic goals for capacity are overly broad and fail to address the specific requirements of individual regions and airspaces. To sum up, although the low-altitude ATM development is progressing, it remains in the early stages to support practical implementation.
@@
Moreover, the deployment of LAWNs necessitates a strategic alignment with both international frameworks and domestic regulations to ensure safety, interoperability, and compliance. A comprehensive review of representative standards relevant to LAWNs is provided to demonstrate how such systems can seamlessly integrate into global and national UAV ecosystems while adhering to regulatory requirements. The ICAO has developed the ATM framework, which outlines essential capabilities for managing UAVs within LAA [67]. This framework emphasizes the need for global harmonization in airspace management and the integration of UAVs into existing ATM systems. Key capabilities within the ATM system include flight authorization, real-time information sharing, conflict management, and automated traffic de-confliction, all of which must be supported by LAWNs to ensure safe and coordinated UAV operations. By doing so, LAWNs can ensure the safe management of UAVs, complementing the existing aviation infrastructure.
@@
In the United States, the Federal Aviation Administration (FAA) introduced its ATM concept of operations [50], envisioning a network of UAV service suppliers that provide services such as digital flight planning, airspace management, and UAV tracking. The FAA’s framework operates through cloud-to-cloud data exchanges, eliminating the need for traditional voice communication with air traffic control. A key feature of the FAA system is Remote ID, which mandates UAVs to broadcast unique identification and location data to enable real-time tracking and compliance with airspace restrictions. LAWNs must integrate with these services, ensuring that UAVs are monitored, controlled, and managed in accordance with FAA regulations. To support studies on future parcel delivery by UAV, the Japan Aerospace Exploration Agency (JAEA) has introduced a dedicated simulation platform for the 2030 timeframe, termed the scalable simulator for knowledge of low-altitude environment (SKALE). This framework emulates BVLoS operations above densely populated urban regions. As illustrated in Fig. 10, a representative delivery scenario is configured with a nominal cruise altitude of 120 m. Analysis based on SKALE indicates that current conflict detection and resolution strategies are not yet sufficient for large-scale, highly automated traffic. In particular, the appropriate setting and modelling of separation minima and flight speeds remain open research problems that require further refinement.
Figure 10: The typical UAV delivery workflow of Ref. [124].
@@
Similarly, the Civil Aviation Administration of China (CAAC) has implemented regulations that follow a risk-based approach to managing LAA [25]. UAVs are categorized based on their weight and operational risk level, with all UAVs, except the smallest, required to be equipped with unique electronic identification and registered under a real-name system. In addition, LAWNs are required to integrate geofencing capabilities and enforce remote flight restrictions according to the operational limitations set by regulatory authorities. The policies within China are actively driving the LAE development, generating significant demand for high-capacity, scalable LAWNs that can support dense UAV applications in terms of drone delivery, eVTOL air taxis, and remote sensing. To meet these requirements, LAWNs must provide ultra-reliable communication and low-latency connectivity to ensure seamless integration and compliance with regulatory standards in an expanding LAA. Toward this end, the 3rd Generation Partnership Project (3GPP) has released several technical reports (TRs) addressing LAWN communications [1]. Notably, TR 36.777 explores the UAV connectivity by leveraging LTE, highlighting key challenges such as interference and mobility. Furthermore, TR 38.811 and TR 38.821 focus on the integration of 5G new radio (NR) within non-terrestrial networks (NTN), which includes satellites and HAPs.
@@ 7.3 Emerging Path Planning Technology
The preceding discussion on low-altitude ATM focuses on macroscopic aspects such as capacity, separation policy, and regulatory coordination. In practice, these system-level concepts are ultimately realized through the individual trajectories flown by each aerial vehicle. Therefore, path planning forms the microscopic counterpart of ATM and directly determines whether a low-altitude airspace can accommodate dense operations without compromising safety or efficiency. A well-designed trajectory must remain compatible with corridor structures, restricted zones, and dynamic flow constraints, while obeying platform limits on energy, payload, and maneuverability. In the context of LAWNs, where numerous heterogeneous UAVs operate in close proximity, deficiencies in path design manifest as conflict hot spots, unnecessary detours, and excessive delay. From a design perspective, most emerging path planning technologies are driven by two primary objectives, namely, trajectory safety and trajectory efficiency [4]. Specifically, safety-oriented planning emphasizes collision and obstacle avoidance, while compliance with no-fly regions and separation minima. On the other hand, efficiency-oriented planning seeks to reduce mission cost, for example, by shortening flight time and lowering energy consumption through multiobjective criteria. In many recent studies, safety is enforced through hard constraints, whereas efficiency guides the selection of a preferred solution within the feasible set.
@@
By contrast, efficiency-oriented path planning methods build on this safety foundation and place additional emphasis on operating cost and computational complexity. For instance, the work in [181] modeled UAV path planning under partial observability as a partially observable Markov decision process and sought trajectories that minimize cumulative operating cost while accounting for nonlinear dynamics and measurement uncertainty. This formulation provides a principled way to balance performance and robustness when the vehicle has only partial knowledge of its environment. Along this line, the work in [38] observed that traditional UAV path planners often ignore mission-specific constraints such as terminal heading and proposed a bidirectional ant-colony-optimization scheme on a grid-based workspace to address this limitation. Ant colonies are initialized near the start and goal configurations and expand simultaneously along predetermined directions, after which a dedicated path-selection mechanism combines and filters candidate routes. Improved pheromone-update rules and successor-selection strategies were introduced to accelerate convergence and reduce the risk of local optima. From a cooperative perspective, the authors in [20] developed a distributed controller for a fleet of UAVs under a multi-agent-systems framework, where online path planning is formulated as a particle-swarm-optimization (PSO) problem solved independently on each vehicle. By minimizing a suitably designed cost function, each local PSO instance determines a trajectory that contributes to formation keeping, target tracking, and collision avoidance, and experiments with fixed and moving targets, external disturbances, and agent loss confirmed that the distributed strategy can maintain coherent fleet behavior and safety without centralized coordination. Overall, emerging path planning technologies complement low-altitude ATM by bridging system-level policies with executable trajectories. Safety-oriented configuration-space methods and cooperative decision frameworks provide reliable guarantees in structured and moderately dynamic airspace, whereas efficiency-oriented approaches enhance scalability and performance in large-scale, high-density LAWNs. The combined use of these approaches is expected to be essential for supporting dense, task-driven operations while maintaining both operational safety and mission effectiveness.
@@ 9 Conclusion
Acknowledgment
This work is supported in part by National Natural Science Foundation of China under Grant 62501600 and 62471208, in part by Guangdong Provincial Natural Science Foundation under Grant 2024A151510098, in part by Shenzhen Science and Technology Program under Grant JCYJ20240813094627037.